OpenStudy (anonymous):

(z-(1/z))/(1-(1/z)) ill rewright it after I post so it is easier to read

OpenStudy (anonymous):



OpenStudy (zehanz):

\[\frac{ z-\frac{ 1 }{ z } }{ 1-\frac{ 1 }{ z } }=\frac{ \frac{ z }{ z }-\frac{ 1 }{ z } }{ \frac{ z^2 }{ z }-\frac{ 1 }{ z } }=\frac{ \frac{ z^2-1 }{ z } }{ \frac{ z-1 }{ z } }\]Does this help? There is more to be done...

OpenStudy (zehanz):

Typo: the z² in the denominator of the middle term must move up 2 positions ;). In the last term everything is fine..

OpenStudy (anonymous):

is the answer z+1?

OpenStudy (anonymous):

zehanz

OpenStudy (zehanz):

it is! you 'll get this:\[\frac{ z^2-1 }{ z }*\frac{ z }{ z-1 }=\frac{ z(z-1)(z+1) }{ z(z-1) }\]So only z+1 is left.

OpenStudy (anonymous):

Thank you

OpenStudy (zehanz):

yw!